# Leveraging parallelizability and channel structure in THz-band, Tbps channel-code decoding

### v1

@@ IV-B Parallelism via short codes
The architecture of the coding scheme plays a crucial role in decoding parallelism, with inherently parallel decoders being preferred in this context. However, strategies for parallel decoding can significantly impact throughput, latency, and energy efficiency. Significant advancements have been made in THz/Tbps coding schemes under $28\,\mathrm{nm}$ technology. LDPC codes have achieved an energy efficiency of $6\,\mathrm{pJ/bit}$ and a throughput of $160\,\mathrm{Gbps}$. Polar codes have reached a throughput of $620\,\mathrm{Gbps}$ with an energy efficiency of $4.6\,\mathrm{pJ/bit}$. However, turbo codes only achieved a throughput of $48\,\mathrm{Gbps}$ and an area efficiency of $5.33\,\mathrm{W/mm^{2}}$ [7].
@@ V Simulation results
Fig. 1a illustrates the performance within a frequency-selective THz channel under ZF detection, with small-scale fading modeled using an $\alpha$-$\mu$ distribution. Simulations compare SCL and GRAND decoding of two polar-code configurations with the following effective rates: $K/N=116/128$ and $K/N=57/64$, both with a cyclic redundancy check (CRC) of 11 bits. Notably, significant performance discrepancies exist among hard, PSI-based, and soft detection. The gap between hard and soft detection exceeds 12 dB at a block error rate (BLER) of $10^{-3}$. However, the gap between PSI-based and soft detection is approximately 4 dB at the same BLER level, substantiating the efficacy of PSI in designing low-complexity non-iterative THz detectors and decoders. Additionally, GRAND exhibits superior performance, outperforming SCL decoding by 2 dB. This superiority is attributed to GRAND’s capability to consider both CRC and polar code redundancies simultaneously.

### v2

@@ IV-B Parallelism via short codes
The architecture of the coding scheme plays a crucial role in decoding parallelism, with inherently parallel decoders being preferred in this context. However, strategies for parallel decoding can significantly impact throughput, latency, and energy efficiency. Significant advancements have been made in THz/Tbps coding schemes under $28\,\mathrm{nm}$ technology. LDPC codes have achieved an energy efficiency of $6\,\mathrm{pJ/bit}$ and a throughput of $160\,\mathrm{Gbps}$. Polar codes have reached a throughput of $620\,\mathrm{Gbps}$ with an energy efficiency of $4.6\,\mathrm{pJ/bit}$. However, turbo codes only achieved a throughput of $48\,\mathrm{Gbps}$ and an area efficiency of $5.33\,\mathrm{W/mm^{2}}$ [7]. Soft-detection ordered reliability bits GRAND (ORBGRAND) has also demonstrated an energy efficiency of $0.76\,\mathrm{pJ/bit}$ and power consumption of $4.9\,\mathrm{mW}$ at $10^{-7}$ frame-error-rate [22], achieving an area efficiency of $16.3\,\mathrm{Gbps/mm^{2}}$ in $40\,\mathrm{nm}$ CMOS.
@@ V Simulation results
Fig. 1a illustrates the performance within a frequency-selective THz channel under ZF detection, with small-scale fading modeled using an $\alpha$-$\mu$ distribution. Simulations compare SCL and soft GRAND (ORBGRAND) [24] decoding of two polar-code configurations with the following effective rates: $K/N=116/128$ and $K/N=57/64$, both with a cyclic redundancy check (CRC) of 11 bits. Notably, significant performance discrepancies exist among hard, PSI-based, and soft detection. The gap between hard and soft detection exceeds 12 dB at a block error rate (BLER) of $10^{-3}$. However, the gap between PSI-based and soft detection is approximately 4 dB at the same BLER level, substantiating the efficacy of PSI in designing low-complexity non-iterative THz detectors and decoders. Additionally, ORBGRAND exhibits superior performance, outperforming SCL decoding by 2 dB. This superiority is attributed to GRAND’s capability to consider both CRC and polar code redundancies simultaneously. Further gains in soft GRAND can be achieved by utilizing the optimal, though more complex, SGRAND [25].